# CDQuant: Accurate Post-training Weight Quantization of Large Pre-trained Models using Greedy Coordinate Descent

### v2

@@ Abstract
Large language models (LLMs) Work done when the authors were at Google Research. have recently demonstrated remarkable performance across diverse language tasks. But their deployment is often constrained by their substantial computational and storage requirements. Quantization has emerged as a key technique for addressing this challenge, enabling the compression of large models with minimal impact on performance. The recent GPTQ algorithm, a post-training quantization (PTQ) method, has proven highly effective for compressing LLMs, sparking a wave of research that leverages GPTQ as a core component. Recognizing the pivotal role of GPTQ in the PTQ landscape, we introduce CDQuant, a simple and scalable alternative to GPTQ with improved performance. CDQuant uses coordinate descent to minimize the layer-wise reconstruction loss to achieve high-quality quantized weights. Our algorithm is easy to implement and scales efficiently to models with hundreds of billions of parameters. Through extensive evaluation on the PaLM2 model family, we demonstrate that CDQuant consistently outperforms GPTQ across diverse model sizes and quantization levels. In particular, for INT2 quantization of PaLM2-Otter, CDQuant achieves a $10\%$ reduction in perplexity compared to GPTQ.
@@ 1 Introduction
Given the central role GPTQ plays in the landscape of PTQ methods for LLMs, improving its quality directly translates to better quantization across numerous techniques that build upon it. Consequently, in this work, we revisit the core optimization problem addressed by GPTQ and introduce a novel coordinate descent based algorithm, CDQuant, to minimize the objective. CDQuant is an iterative optimization technique that greedily selects coordinates to descend along in each iteration. This approach contrasts with GPTQ, which cycles through the coordinates only once, in a predetermined order, often leading to suboptimal solutions for the layer-wise objective (see Section 2 for more details). CDQuant is simple to implement and scales effectively to models with billions of parameters. We further extend CDQuant  to group/sub-channel quantization, a technique gaining popularity of late (Dettmers et al., 2021; Dettmers and Zettlemoyer, 2023). Extensive experiments on PaLM 2 (Anil et al., 2023) demonstrate that CDQuant consistently outperforms GPTQ across various model sizes and quantization precision levels. For example, for INT2 quantization of PaLM2-Otter, CDQuant achieves a $10\%$ reduction in perplexity compared to GPTQ.
@@ 2 Related Work
Optimal Brain Quantization (OBQ) (Frantar and Alistarh, 2022), a post-training quantization (PTQ) technique, draws inspiration from the Optimal Brain Surgeon (OBS) framework proposed by LeCun et al. (1989). Frantar and Alistarh (2022) introduced exact yet efficient algorithms for implementing the OBS greedy strategy. However, OBQ remains computationally expensive and struggles to scale effectively to large language models (LLMs). Frantar et al. (2022) subsequently introduced GPTQ, which employs heuristics to accelerate the OBQ algorithm. Specifically, GPTQ updates coordinates once in a cyclic manner rather than the greedy approach used in OBQ. This modification significantly speeds up the algorithm but comes at the cost of reduced performance. In our work, we address this performance drop by proposing greedy coordinate descent algorithms that are both straightforward and easy to implement.
@@
Recent works have sought to improve upon GPTQ. One popular strategy here is to identify and isolate problematic weights before applying GPTQ. For example, SpQR (Dettmers et al., 2023) isolates outlier weights and keeps them at full precision, quantizing only the remaining values using GPTQ. Similarly, OWQ (Lee et al., 2023) identifies weights that are highly sensitive to small perturbations and excludes them from the GPTQ quantization process. Another promising direction involves transforming the weights into a different basis before quantization. Both QuIP (Chee et al., 2024), and FrameQuant (Adepu et al., 2024) take this route and perform quantization in the transformed space using GPTQ. In all these approaches, we could potentially substitute GPTQ with CDQuant and expect to see further improvements in performance. Other promising methods such as AWQ (Lin et al., 2023), AffineQuant (Ma et al., 2024), suppress the effect of outliers in input activations by adjusting their scale and transferring it to weights. They then perform the standard MinMax quantization on the rescaled weights. One could replace MinMax with GPTQ, CDQuant to improve the performance of these techniques (see Table 3 for a comparison of CDQuant with AWQ).
@@
While working on our paper, we came across QuantEase (Behdin et al., 2023), a parallel research effort sharing similar goals as ours for improving GPTQ. While both methods leverage the concept of coordinate descent, QuantEase adopts a cyclic approach for weight updates, whereas our work employs a greedy strategy. Although our experiments (Appendix F) indicate comparable performance between the two methods, our work extends beyond QuantEase by introducing specialized algorithms for group/sub-channel quantization, not just full-channel quantization. Additionally, we develop novel block coordinate descent algorithms that further improve the performance. That being said, both these works (QuantEase and ours) collectively highlight the potential of coordinate descent algorithms in outperforming GPTQ.
Weight+Activation Quantization.
@@
SmoothQuant (Xiao et al., 2023), OS+(Wei et al., 2023, 2022) have a similar flavour as AWQ, but quantize both weights and activations after scaling them appropriately. OmniQuant (Shao et al., 2023) performs quantization of the entire transformer block in a single shot. This encompasses both activation and weight quantization. Furthermore, it subsumes SmoothQuant, OS+ by using both feature scaling and outlier suppression. QLLM (Liu et al., 2023) tackles the issue of outliers in the activations by splitting the outlier features into multiple sub-channels and then recombining them, effectively reducing their influence. QLLM also incorporates a fine-tuning step at the end, introducing low-rank weights into each layer of the LLM. LLM.int8() (Dettmers et al., 2022) quantizes both acativations and weights to $8$-bits and also identifies outliers and stores them in full precision.
Efficient End-to-End Quantization.
@@
To recover the drop in performance from quantization Chai et al. (2023); Dettmers et al. (2024); Hu et al. (2021) perform low-rank parameter-efficient fine-tuning.
@@ 3 CDQuant
For a fixed $(a,b)$, this problem is called Integer Linear Regression problem. It turns out, finding optimal solutions to this problem is NP-hard (Chrétien and Corset, 2009; Park and Boyd, 2018). So, several works have designed heuristics to solve this problem (Nagel et al., 2020; Li et al., 2021; Frantar and Alistarh, 2022; Frantar et al., 2022; Hubara et al., 2021). Within the context of LLMs, GPTQ is perhaps the most popular among these techniques, as it scales efficiently to models with billions of parameters (Frantar et al., 2022). In this work, we aim to improve upon GPTQ by designing better heuristics, while maintaining its scalability aspect. To this end, we rely on performing coordinate descent on objective (1), which we describe below.
@@ 3.1 Greedy Coordinate Descent
$\min_{i,r}{\mathcal{L}}(\mathbf{q}+(r-\mathbf{q}_{i})\mathbf{e}_{i})-{\mathcal{L}}(\mathbf{q}),$
where $\mathbf{e}_{i}$ is the standard basis vector with $1$ in $i$ position and 00 everywhere else. Luckily for us, this can be implemented extremely efficiently as we have analytical expressions for the objective. In particular, one can easily show that
${\mathcal{L}}(\mathbf{q}+(r-\mathbf{q}_{i})\mathbf{e}_{i})-{\mathcal{L}}(\mathbf{q})=(r-\mathbf{q}_{i})^{2}H_{i,i}+(r-\mathbf{q}_{i})\mathbf{g}_{i},$
where $H,\mathbf{g}$ are the Hessian and gradient of ${\mathcal{L}}$ evaluated at $\mathbf{q}$. Algorithm 1 describes this procedure. Observe that line 5 is the most expensive step of the algorithm. It requires finding the minimum value among $d_{\text{in}}\times 2^{c}$ possibilities, a task that is well-suited for parallelization on GPUs.
Extension to Block Coordinate Descent.
A natural extension to Algorithm 1 is block coordinate descent (BCD), where multiple coordinates are updated simultaneously in each iteration. While a greedy approach to BCD could, in principle, optimize the objective much better, the computational cost becomes prohibitive. Specifically, updating $k$ coordinates at a time necessitates evaluating $(d_{\text{in}}\times 2^{c})^{k}$ possible combinations of coordinates and their corresponding values. To overcome this, we propose a randomized BCD strategy (Algorithm 2), where we randomly partition the coordinates into $d_{\text{in}}/k$ blocks, and only search over these blocks and their corresponding values. This significantly reduces the search space to a more manageable $d_{\text{in}}/k\times 2^{kc}$ possibilities, making the algorithm practical for larger models. In our experiments, we primarily set $k=2$ and work with $c\in\{2,3,4\}$.
Initializing $a,b,\mathbf{q}$.
To initialize $a,b,\mathbf{q}$ in Algorithms 12, we introduce a technique called Optimal Weight Clipping (OWC), which draws inspiration from the Learnable Weight Clipping (LWC) mechanism used in OmniQuant (Shao et al., 2023). In OWC, we quantize weight $\mathbf{w}$ as follows
@@
Note that while not explicitly stated, both $\mathbf{w},a$ in the above objective are implicitly dependent on $\gamma$. This optimization can be efficiently solved using a simple grid search. In contrast, LWC (Shao et al., 2023) optimizes a different objective function, focusing on end-to-end quantization of entire transformer block using gradient-based techniques. It is worth noting that setting $\gamma=1$ in OWC recovers the widely used MinMax quantization scheme, which is used in many existing quantization methods, including GPTQ (Frantar et al., 2022), SmoothQuant‘(Xiao et al., 2023). In our experiments, we observed that OWC provides a much better initialization compared to MinMax quantization, for both GPTQ and CDQuant.
@@ 3.2 Extension to Sub-channel Quantization
In this section, we consider sub-channel (or group) quantization, which is a more fine-grained quantization that divides the weight vector $\mathbf{w}$ into multiple groups and assigns a quantization scale to each group (Dettmers et al., 2021; Dettmers and Zettlemoyer, 2023). Letting $g$ be the group size, we divide weight $\mathbf{w}$ into $d_{\text{in}}/g$ groups $\{\mathbf{w}^{(0)},\mathbf{w}^{(1)},\dots\mathbf{w}^{(d_{\text{in}}/g-1)}\}$ each of size $g$, and quantize $\mathbf{w}^{(i)}$ as $a^{(i)}\mathbf{q}^{(i)}+b^{(i)}$. To learn the optimal parameters for this sub-channel quantization, we solve the following optimization problem:
@@
Here, $X^{(i)}$ represents the columns of $X$ corresponding to the indices within group $i$. To solve this optimization problem, we employ a coordinate descent approach, similar to Algorithms 1 and 2 described earlier. That is, given initial values for the scaling and bias parameters, we iteratively optimize the quantized representation $\mathbf{q}$ using coordinate descent. Due to space constraints, we present the resulting algorithms in Appendix C (see Algorithms 45).
@@
We use greedy coordinate descent to optimize Equation (4). In each iteration, we update the $\gamma^{(i)}$ that leads to biggest drop in loss. This procedure, which we call OWC-CD, is described in Algorithm 3.
@@
$\Delta(i,\beta)\leftarrow\mathbf{w}^{(i)}-a^{(i)}(\beta)\mathbf{q}^{(i)}(\beta)-b^{(i)}(\beta)$
.    where $a^{(i)}(\beta),b^{(i)}(\beta),\mathbf{q}^{(i)}(\beta)$ are as defined in Equation (2).

### v3

# CDQuant: Greedy Coordinate Descent for Accurate LLM Quantization

@@ Abstract
Large language models (LLMs) have recently demonstrated remarkable performance across diverse language tasks. But their deployment is often constrained by their substantial computational and storage requirements. Quantization has emerged as a key technique for addressing this challenge, enabling the compression of large models with minimal impact on performance. The recent GPTQ algorithm, a post-training quantization (PTQ) method, has proven highly effective for compressing LLMs, sparking a wave of research that leverages GPTQ as a core component. Recognizing the pivotal role of GPTQ in the PTQ landscape, we introduce CDQuant, a simple and scalable alternative to GPTQ with improved performance. CDQuant uses greedy coordinate descent to minimize the layer-wise reconstruction loss to achieve high-quality quantized weights. Our algorithm is easy to implement and scales efficiently to models with hundreds of billions of parameters. We perform extensive evaluation on Gemma, and PaLM2 model families, and demonstrate that CDQuant consistently outperforms GPTQ in 2-4 bit weight quantization. Moreover, CDQuant improves the performance of state-of-the-art PTQ techniques such as QuIP and FrameQuant when used as a replacement for their GPTQ component, resulting in further gains in quality.
@@ 1 Introduction
Given the central role GPTQ plays in the landscape of PTQ methods for LLMs, improving its quality directly translates to better quantization across numerous techniques that build upon it. Consequently, in this work, we revisit the core optimization problem addressed by GPTQ and introduce a novel coordinate descent based algorithm, CDQuant, to minimize the objective. CDQuant is an iterative optimization technique that greedily selects coordinates to descend along in each iteration. This approach contrasts with GPTQ, which cycles through the coordinates only once, in a predetermined order, often leading to suboptimal solutions for the layer-wise objective (see Section 2 for more details). CDQuant is simple to implement and scales effectively to models with billions of parameters. We further extend CDQuant  to group/sub-channel quantization, a technique gaining popularity of late (Dettmers et al., 2021; Dettmers & Zettlemoyer, 2023).
Extensive experiments on Gemma (Mesnard et al., 2024; Gemma-Team, 2024), and PaLM2 (Anil et al., 2023) model families demonstrate that CDQuant consistently outperforms GPTQ across various model sizes and quantization precision levels (2-4 bits). Notably, for INT2 quantization of PaLM2-Otter, CDQuant achieves a $10\%$ perplexity reduction compared to GPTQ. Furthermore, integrating CDQuant with QuIP (Chee et al., 2024) and FrameQuant (Adepu et al., 2024) improves their performance by $\sim 5\%$ for INT2 quantization of the Gemma-2 27B model, showcasing its potential as a drop-in replacement for GPTQ in existing PTQ techniques.
@@ 2 Related Work
GPTQ. Inspired by the Optimal Brain Surgeon (OBS) framework for pruning (LeCun et al., 1989), Optimal Brain Quantization (OBQ) (Frantar & Alistarh, 2022) emerged as a promising post-training quantization (PTQ) technique. However, OBQ remains computationally expensive and struggles to scale effectively to large language models (LLMs). This is because of a complicated step involved in OBQ which requires updating all the unquantized coordinates in each step (see Equation 2 in Frantar et al. (2022)). Implementing this step in GPUs is extremely slow because of the sheer number of gathers/scatters that need to be performed at each step. Frantar et al. (2022) subsequently introduced GPTQ, which employs heuristics to accelerate the OBQ algorithm. Specifically, GPTQ updates coordinates once in a cyclic manner rather than the greedy approach used in OBQ. This modification significantly speeds up the algorithm but comes at the cost of reduced quality. In our work, we address this performance drop by proposing greedy coordinate descent algorithms that are both straightforward and easy to implement.
Other Weight-only Quantization Techniques. Recent works on post-training quantization have sought to improve upon GPTQ. One popular strategy here is to identify and isolate problematic weights before applying GPTQ. For example, SpQR (Dettmers et al., 2023) isolates outlier weights and keeps them at full precision, quantizing only the remaining values using GPTQ. Similarly, OWQ (Lee et al., 2023) identifies weights that are highly sensitive to small perturbations and excludes them from the GPTQ quantization process. JSQ (Guo et al., 2024) clips activations (where the clipping thresholds are searched through simulated annealing), following which a combination of activation range and weight saliency are used to sparsify the weight matrix before quantizing it with GPTQ. SqueezeLLM (Kim et al., 2024) also preserves outliers in floating point, however, employs non-uniform quantization by performing K-means clustering on the final end-to-end loss (where the gradient is approximated with zero,and the Hessain with the Fisher information matrix). Any-Precision LLM (Park et al., 2024) improves SqueezeLLM by allowing several nested precisions through a control over the number of cluster centers. However, both these approaches involve storing a codebook, hence making hardware and software compatibility tricky.
@@
Another promising direction involves transforming the weights into a different basis before quantization. QuIP (Chee et al., 2024), and FrameQuant (Adepu et al., 2024), SpinQuant (Liu et al., 2024), QuaRot (Ashkboos et al., 2024) take this route and perform quantization in the transformed space using GPTQ. In all these approaches, we could potentially substitute GPTQ with CDQuant and expect to see further improvements in performance (see Section 4 for empirical evidence). Other promising methods such as AWQ (Lin et al., 2023), AffineQuant (Ma et al., 2024), suppress the effect of outliers in input activations by adjusting their scale and transferring it to weights. They then perform the standard MinMax quantization on the rescaled weights. One could replace MinMax with GPTQ, CDQuant to improve the performance of these techniques (see Table 4 for a comparison of CDQuant with AWQ).
@@
While working on our paper, we came across QuantEase (Behdin et al., 2023), a parallel research effort sharing similar goals as ours for improving GPTQ. While both methods leverage the concept of coordinate descent, QuantEase adopts a cyclic approach for weight updates, whereas our work employs a greedy strategy. Although our experiments (Appendix D) indicate comparable performance between the two methods, our work extends beyond QuantEase by introducing specialized algorithms for group/sub-channel quantization, not just full-channel quantization. Additionally, we develop novel block coordinate descent algorithms that further improve the performance. That being said, both these works (QuantEase and ours) collectively highlight the potential of coordinate descent algorithms in outperforming GPTQ.
Vector Quantization. AQLM (Egiazarian et al., 2024) generalizes Additive Quantization to LLMs by introducing input-adaptiveness (through calibration data) and jointly optimizing the codebooks across transformer blocks. QuIP# (Tseng et al., 2024) improves QuIP’s incoherence processing through random Hadamard matrices, introduces vector quantization, and adds an additional layer of finetuning to improve model quality. However, in this work, our focus is methods that improve scalar quantization.
@@
Weight+Activation Quantization. SmoothQuant (Xiao et al., 2023), OS+(Wei et al., 2023; 2022) have a similar flavour as AWQ, but quantize both weights and activations after scaling them appropriately. OmniQuant (Shao et al., 2023) performs quantization of the entire transformer block in a single shot. This encompasses both activation and weight quantization. Furthermore, it subsumes SmoothQuant, OS+ by using both feature scaling and outlier suppression. QLLM (Liu et al., 2023) tackles the issue of outliers in the activations by splitting the outlier features into multiple sub-channels and then recombining them, effectively reducing their influence. QLLM also incorporates a fine-tuning step at the end, introducing low-rank weights into each layer of the LLM. LLM.int8() (Dettmers et al., 2022) quantizes both acativations and weights to $8$-bits and also identifies outliers and stores them in full precision. LQER (Zhang et al., 2024) approximates the weight quantization error with the product of two low rank matrices.
@@
Efficient End-to-End Quantization. To recover the drop in performance from quantization Chai et al. (2023); Dettmers et al. (2024); Hu et al. (2021) perform low-rank parameter-efficient fine-tuning.
@@ 3 CDQuant
For a fixed $(a,b)$, this problem is called Integer Linear Regression problem. It turns out, finding optimal solutions to this problem is NP-hard (Chrétien & Corset, 2009; Park & Boyd, 2018). So, several works have designed heuristics to solve this problem (Nagel et al., 2020; Li et al., 2021; Frantar & Alistarh, 2022; Frantar et al., 2022; Hubara et al., 2021). Within the context of LLMs, GPTQ is perhaps the most popular among these techniques, as it scales efficiently to models with billions of parameters (Frantar et al., 2022). In this work, we aim to improve upon GPTQ by designing better heuristics, while maintaining its scalability aspect. To this end, we rely on performing coordinate descent on objective (1), which we describe below.
@@ 3.1 Greedy Coordinate Descent
3:Compute gradient $\mathbf{g}$ at $\mathbf{q}_{0}$ as: $\mathbf{g}\leftarrow 2H(\mathbf{q}_{0}-a^{-1}(\mathbf{w}-b))$
@@
$\min_{i,r}{\mathcal{L}}(\mathbf{q}+(r-q_{i})\mathbf{e}_{i})-{\mathcal{L}}(\mathbf{q}),$
where $q_{i}$ is the $i^{th}$ element of $\mathbf{q}$, and $\mathbf{e}_{i}$ is the standard basis vector with $1$ in $i$ position and 00 everywhere else. Luckily for us, this can be implemented extremely efficiently as we have analytical expressions for the objective. In particular, one can easily show that
${\mathcal{L}}(\mathbf{q}+(r-q_{i})\mathbf{e}_{i})-{\mathcal{L}}(\mathbf{q})=(r-q_{i})^{2}H_{i,i}+(r-q_{i})\mathbf{g}_{i},$
where $H,\mathbf{g}$ are the Hessian and gradient of ${\mathcal{L}}$ evaluated at $\mathbf{q}$. This follows from the fact that ${\mathcal{L}}$ is a quadratic function in $\mathbf{q}$. Algorithm 1 describes this procedure (note that line 5 is the key step which identifies the descent direction).
Remark 1 (Comparison with OBQ).
Similar to our technique, OBQ also updates its coodinates in a greedy manner. But it differs from our algorithm in one crucial step: it updates all the unquantized coordinates in each step (see Equation 2 in Frantar et al. (2022)). Implementing this step on accelerators such as GPUs/TPUs is extremely slow because of the sheer number of gathers/scatters that need to be performed at each step. In contrast, our algorithm only requires updating one coordinate at a time.
Extension to Block Coordinate Descent. A natural extension to Algorithm 1 is block coordinate descent (BCD), where multiple coordinates are updated simultaneously in each iteration. While a greedy approach to BCD could, in principle, optimize the objective much better, the computational cost becomes prohibitive. Specifically, updating $k$ coordinates at a time necessitates evaluating $(d_{\text{in}}\times 2^{c})^{k}$ possible combinations of coordinates and their corresponding values. To address this, we propose a randomized BCD strategy (see Algorithm 2), which partitions the coordinates into random $d_{\text{in}}/k$ blocks and searches only over these blocks. This significantly reduces the search space to a more manageable $d_{\text{in}}/k\times 2^{kc}$ possibilities, making the algorithm practical for larger models. In our experiments, we primarily use $k=2$ and $c\in\{2,3,4\}$. Note that Algorithm 2 with $k=1$ recovers Algorithm 1.
Initializing $a,b,\mathbf{q}$. To initialize $a,b,\mathbf{q}$ in Algorithms 12, we introduce a technique called Optimal Weight Clipping (OWC), which draws inspiration from the Learnable Weight Clipping (LWC) mechanism used in OmniQuant (Shao et al., 2023). In OWC, we quantize weight $\mathbf{w}$ as follows
@@
Note that while not explicitly stated, both $\mathbf{q},a$ in the above objective are implicitly dependent on $\gamma$. This optimization can be efficiently solved using a simple grid search. In contrast, LWC (Shao et al., 2023) optimizes a different objective function, focusing on end-to-end quantization of entire transformer block using gradient-based techniques. It is worth noting that setting $\gamma=1$ in OWC recovers the widely used MinMax quantization scheme, which is used in many existing quantization methods, including GPTQ (Frantar et al., 2022), SmoothQuant(Xiao et al., 2023). However, MinMax quantization is susceptible to outlier weights, and a smaller $\gamma$ often yields superior results. In our experiments, we observed that OWC provides a much better initialization compared to MinMax quantization, for both GPTQ and CDQuant, suggesting its broader applicability.
@@
$i^{*},\mathbf{r}^{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}i\in\{0,1,\dots d_{\text{in}}/k-1\},\\ \mathbf{r}\in\{0,1,\dots 2^{c}-1\}^{k}\end{subarray}}(\mathbf{r}-\mathbf{q}_{t-1,i})^{T}H_{i,i}(\mathbf{r}-\mathbf{q}_{t-1,i})+(\mathbf{r}-\mathbf{q}_{t-1,i})^{T}\mathbf{g}_{i},$
@@
$\mathbf{g}\leftarrow\mathbf{g}+2H_{i^{*},\cdot}(\mathbf{r}^{*}-\mathbf{q}_{t-1,i^{*}}),$
@@
$\mathbf{q}_{t-1,i^{*}}\leftarrow\mathbf{r}^{*},\quad\mathbf{q}_{t}\leftarrow\mathbf{q}_{t-1},$
@@ 3.2 Extension to Sub-channel Quantization
In this section, we consider sub-channel (or group) quantization, which is a more fine-grained quantization that divides the weight vector $\mathbf{w}$ into multiple groups and assigns a quantization scale to each group (Dettmers et al., 2021; Dettmers & Zettlemoyer, 2023). Letting $g$ be the group size, we divide weight $\mathbf{w}$ into $d_{\text{in}}/g$ groups $\{\mathbf{w}^{(0)},\mathbf{w}^{(1)},\dots\mathbf{w}^{(d_{\text{in}}/g-1)}\}$ each of size $g$, and quantize $\mathbf{w}^{(i)}$ as $a^{(i)}\mathbf{q}^{(i)}+b^{(i)}$. To learn the optimal parameters for this sub-channel quantization, we solve the following optimization problem:
@@
Here, $X^{(i)}$ represents the columns of $X$ corresponding to the indices within group $i$. To solve this optimization problem, we employ a coordinate descent approach, similar to Algorithms 1 and 2 described earlier. That is, given initial values for the scaling and bias parameters, we iteratively optimize the quantized representation $\mathbf{q}$ using coordinate descent. Due to space constraints, we present the resulting algorithms in Appendix E (see Algorithms 45).
@@
While not explicitly stated, both $\mathbf{q}^{(i)},a^{(i)}$ implicitly depend on the clipping strength $\gamma^{(i)}.$ We use greedy coordinate descent to optimize Equation (4). In each iteration, we update the $\gamma^{(i)}$ that leads to biggest drop in loss (lines 10-14 of Algorithm 3). This procedure, which we call OWC-CD, is described in Algorithm 3.
@@
$\Delta(i,\beta)\leftarrow\mathbf{w}^{(i)}-a^{(i)}(\beta)\mathbf{q}^{(i)}(\beta)-b^{(i)}(\beta),$
   where $a^{(i)}(\beta),b^{(i)}(\beta),\mathbf{q}^{(i)}(\beta)$ are as defined in Equation (2).